Circular No. 5625

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1992az IN NGC 818 C. Pollas, Observatoire de la Côte d’Azur, reports his discovery of a

supernova (mag ∼ 18) in the spiral galaxy NGC 818 = UGC 1633. The new object is located at α = 2^{h}05^{m}45^{s}.04, δ = +38^{o}32′10′′.0 (equinox 1950.0), which is 28′′.5 east and 16′′.2 south of the galaxy’s center (end figures 42^{s}.61, 26′′.2). The supernova was found on a Tech Pan film obtained by Pollas (assisted by J.-B. Emond) on Sept. 27.04 UT and confirmed on an R plate taken Sept. 28.08 (respective limiting magnitudes 21.5 and 19). Nothing is visible on the POSS prints to R ∼ 19 and B ∼ 20. A nearby star of mag ∼ 15 has end figures 45^{s}.58, 38′′.1.

F. Michel and P. Prugniel obtained a CCD image on Sept. 29.06 UT with the 1.2-m Observatoire de Haute-Provence telescope, under nonphotometric conditions, measuring V = 18.85 ± 0.2.

R. Kirshner and S. Gordon, Harvard-Smithsonian Center for Astrophysics, communicate: “Spectra obtained with the Multiple Mirror Telescope (+ Red Channel spectrograph, range 400-800 nm) show that this object is a type-II supernova, with a well-developed P-Cyg profile at Hα. It closely resembles other type-II supernovae at an age of ∼ 1 month. This is a reasonable candidate for a distance determination by the Expanding Photosphere Method, and further observations (especially V, R, and I photometry) will be valuable.”

SUPERNOVA 1992ba IN NGC 2082 R. Evans, Hazelbrook, New South Wales, reports his visual discovery

on Sept. 30.75 UT of a supernova of mag ∼ 14.0, located ∼ 1′ west of the center of NGC 2082 (α = 5^{h}41^{m}.6, δ = −64^{o}20′, equinox 1950.0)

M. M. Phillips, Cerro Tololo Interamerican Observatory, reports: “A spectrum obtained of Evans’ supernova in NGC 2082 shows that this object is a type-II event. The spectrum, which was acquired with the CTIO 4.0-m telescope on Oct. 1.36 UT, displays strong, broad, P-Cyg emission lines of hydrogen atop a blue continuum. The absorption component is noticeably weak at Hα, but quite well developed at Hβ and Hγ. Weaker P-Cyg lines of He I (587.6 and 447.1 nm) are clearly visible as well. Synthetic photometry of a wide-slit (10′′) spectrum yields V = 15.22 and B–V = +0.07. From these observed properties, it appears very likely that this supernova has been caught at an early epoch.”

1992 October 1 Daniel W. E. Green